# Grounded in Literature

### v1

@@ Abstract
As AI tools become increasingly common for research ideation, robust evaluation is critical to ensure the validity and usefulness of generated ideas. We introduce ScholarEval, a retrieval-augmented evaluation framework that assesses research ideas based on two fundamental criteria: soundness—the empirical validity of proposed methods based on existing literature, and contribution—the degree of advancement made by the idea across different dimensions relative to prior research. To evaluate ScholarEval, we introduce ScholarIdeas, the first expert-annotated dataset of multi-domain research ideas and reviews, comprised of 117 ideas across four disciplines: artificial intelligence, neuroscience, biochemistry, and ecology. Our evaluation shows that ScholarEval achieves significantly higher coverage of points mentioned in the human expert annotated rubrics in ScholarIdeas compared to all baselines. Furthermore, ScholarEval is consistently preferred over our strongest baseline o4-mini-deep-research, a reasoning and search-enabled agentic system by OpenAI, in terms of evaluation actionability, depth, and evidence support. Our large-scale user study also shows that ScholarEval significantly outperforms deep research in literature engagement, idea refinement, and usefulness. We openly release our code, dataset, and ScholarEval tool for the community to use and build on. [^1] Code and data can be found at https://github.com/skai-research/ScholarEval
@@ 1 Introduction
Figure 1: Left: Given a research idea ScholarEval generates a literature-grounded evaluation based on soundness and contribution. Right: To evaluate ScholarEval, we measure the degree of coverage of expert-annotated review rubrics in ScholarIdeas. The final coverage score is the average over all rubrics.
@@
Despite this need, automatic research idea evaluation remains an underexplored area. Some works narrowly frame it as a prediction task: deciding which idea among a pair would lead to better results on predefined benchmarks (Wen et al., 2025). Others focus on one-dimensional approaches to idea evaluation (e.g., novelty) (Afzal et al., 2025; Shahid et al., 2025), or target only specific sub-disciplines (e.g., AI) (Si et al., 2025b). Most importantly, these systems either only generate scores or are limited to sparse feedback such as short rationale statements (Feng et al., 2025; Wen et al., 2025; Shahid et al., 2025). However, to create AI co-scientists that can generate and refine research ideas, giving dense, actionable, and multifaceted feedback is crucial (Wu et al., 2023; Cao et al., 2024). To the best of our knowledge, no existing work addresses the challenge of comprehensive research idea evaluation across disciplines within a framework that provides detailed and actionable feedback for idea refinement. To address this gap, we introduce ScholarEval (Figure 1), a research idea evaluation system grounded in the most recent literature. ScholarEval evaluates research ideas based on two fundamental criteria: soundness and contribution. (1) Soundness refers to the empirical validity of each proposed method in the research plan, assessed by systematically examining whether similar applications of this method in existing literature have demonstrated success or failure. (2) Contribution refers to the degree of advancement a research idea offers across different dimensions–e.g., its proposed methodology, data, evaluation approaches, and conceptual framework–relative to existing literature. The rationale for dimension-based evaluation is that novelty is multi-faceted by nature, and an idea can be considered novel relative to certain aspects of prior work, rather than being categorically novel or not (Rubaiat et al., 2025; Radensky et al., 2025). Given a research idea detailing the problem, proposed methodology, and planned experiments, ScholarEval employs a multi-stage pipeline that first generates targeted search queries to retrieve a large volume of related literature from Semantic Scholar (Kinney et al., 2025) (Figure 2). It then extracts key information from the retrieved literature to assess the research plan’s soundness and contribution. Finally, it synthesizes detailed feedback supported by relevant citations.
@@
To evaluate ScholarEval, we construct ScholarIdeas, a multi-disciplinary dataset of 117 research ideas and their corresponding reviews validated by subject-matter experts across four disciplines: artificial intelligence, biochemistry, neuroscience, and ecology. As showcased in Figure 1, reviews in ScholarIdeas are composed of multiple rubrics, each focusing on a specific point that the evaluation should address, for 1076 rubrics in total. We develop a multi-faceted automatic evaluation framework to assess ScholarEval against strong baselines, namely state-of-the-art LLMs and deep research systems. Our results show that ScholarEval achieves greater coverage of the expert-annotated review rubrics in ScholarIdeas, significantly outperforming all baselines and surpassing o4-mini-deep-research by over 20% relative improvement. Our results also demonstrate that ScholarEval is consistently preferred over deep research in terms of evidence support, depth, and actionability.
@@
Our human study involving 18 experts and 46 evaluations further supports the real-world usefulness of ScholarEval across our four target disciplines. ScholarEval outperforms the strongest baseline, deep research, by a significant margin in metrics tied to our core contributions: literature engagement, citation use, feedback validity, idea refinement, and focus on relevant evaluation aspects. The expert evaluators also found our system more useful and were eager for its official release.
Our work makes the following major contributions:
@@
  • ScholarEval, a literature-grounded framework that comprehensively evaluates research ideas based on their soundness and contribution with actionable feedback. We openly release the ScholarEval tool.
@@ 2 ScholarEval
Task Formulation. Given a research idea $I$, the task is to find papers $P=\{p_{1},p_{2},p_{3},\ldots\}$ that are highly relevant to $I$ and synthesize their findings in the context of $I$ to generate a comprehensive evaluation $\mathcal{E}=(S,C)$ covering the soundness and contribution of the research idea. The evaluation is accompanied by citations, ensuring that all claims are supported by evidence from the literature.
Overview of ScholarEval. As shown in Figure 2, ScholarEval is composed of two main modules: Soundness and Contribution. We present both their workflows in §2.1 and §2.2 and provide further details in Appendix A.
@@ 2.2 Contribution Evaluation
The initial step of contribution evaluation consists of extracting the dimensions along which the research idea is making contributions to the field. Dimensions represent the facets of the idea’s potential contributions that are specific and comparable across related literature (e.g., system design, data collection, evaluation methodology, etc.). Instead of imposing pre-defined dimensions as in Radensky et al. (2025), we use LLM extraction to ensure flexibility based on the nature of the research idea (Rubaiat et al., 2025). Specifically, given the research idea $I$, we use $\mathcal{M}$ to extract dimensions $D=\{d_{1},d_{2},\ldots,d_{l}\}$. Examples of dimensions include tool or system design, conceptual framework, evaluation methodology, etc. Each $d_{i}$ also includes the reasoning for how the idea makes contributions along that dimension. These statements will provide important context to ground the query generation in the subsequent step.
Paper Discovery.
@@
In this step, we conduct a broad search over the literature to identify relevant papers to compare $I$ against. However, unlike the soundness module, which requires searching paper content for methodological details, contribution evaluation can be performed using only abstracts, since a paper’s main contributions are typically highlighted there. This also allows us to cast a wider net and gather a broad set of related papers. Such breadth is essential for contribution evaluation, as determining truly novel contributions requires an exhaustive view of the literature. The paper discovery thus consists of the following steps: (1) For each extracted dimension, $\mathcal{M}$ generates queries to search for relevant papers using Semantic Scholar paper search and retrieve their abstracts. (2) $\mathcal{M}$ assesses similarity in contributions of each candidate paper abstract relative to the research idea $I$ and assigns a score on a scale from 1 to 5. Papers that are deemed highly relevant (i.e., score $\geq 3$) are then used as seeds for the paper augmentation stage. (3) Paper augmentation leverages the Semantic Scholar Recommendations API to find similar papers, and we additionally extract the publications cited by each seed paper. (4) This augmented list of candidate papers then undergoes another stage of relevance assessment by $\mathcal{M}$. However, due to the typically large volume of papers at this stage, we first filter this list to the top $n$ papers based on semantic embedding [^2] We use Titan Text Embedding v2 (Amazon Web Services, 2025). similarity between abstracts and $I$ before forwarding to $\mathcal{M}$.
Pairwise Comparison.
@@
Once the final set of papers $P_{D}=\{p_{1},p_{2},\ldots,p_{m}\}$ is identified, we conduct a series of pairwise comparisons to uncover how these papers’ contributions compare to those proposed in $I$. Specifically, we prompt $\mathcal{M}$ to compare each paper’s abstract to $I$ along each dimension $d_{i}\in D$. This comparison produces a granular view of the areas in which $I$ is making novel advancements and those in which it is lacking in novelty, compared to existing work.
Contribution Review Synthesis.
@@
The final stage of contribution evaluation consists of synthesizing dimension-level contribution assessments. For each $d_{i}\in D$, $\mathcal{M}$ uses the results of the pairwise comparisons to synthesize an evaluation composed of three sections: (1) Strengths: the novel contributions that $I$ makes along this dimension compared to prior work. (2) Weaknesses: the areas in which the contributions of $I$ are lacking or limited compared to prior work. (3) Suggestions: actionable recommendations to improve the novelty of $I$ along this dimension.
@@ 3 ScholarIdeas: A dataset for research idea evaluation
To this end, we curate ScholarIdeas, a dataset containing research ideas and reviews covering four disciplines: artificial intelligence, neuroscience, biochemistry, and ecology. We employ a semi-automatic pipeline to construct this dataset, where each example is validated by subject-matter experts (§3.1). Our primary motivation for developing ScholarEval is to evaluate AI-generated ideas. However, collecting detailed expert reviews for AI-generated ideas at scale is prohibitively expensive and practically infeasible. Our dataset of existing human-written ideas with expert reviews serves as a reliable proxy.
The evaluation of long-form responses, such as the ones generated by ScholarEval, is also inherently challenging. While other scientific tasks admit easily verifiable success criteria (e.g., code generation (Jansen et al., 2025; Chen et al., 2025; Li et al., 2025b)), such evaluation metrics are not readily available for our use case. We develop a multi-faceted automatic evaluation pipeline detailed in §3.2. We further corroborate our automatic evaluation with a large-scale human evaluation described in §5.
@@ 3.1 Data Curation and Annotation
Data selection. We manually select papers and reviews from two sources: OpenReview (ICLR 2025) for AI-related research ideas and eLife for life sciences. Both sources include high-quality reviews by multiple reviewers for all submitted papers, in contrast to many other sources that do not openly release reviews or restrict them to those of accepted manuscripts. To ensure the suitability of the reviews to be used as ground-truth, we only select papers satisfying the following criteria: (1) The paper must be reviewed by at least two reviewers. (2) All reviews of the paper are in agreement and offer a general consensus about the quality of the work. (3) The reviews mention criticism about the underlying research idea and not exclusively about obtained results or other details known post-execution. For each of the sources we use, we only retrieve the first version of the submission prior to any revisions, along with the first round of reviews. We also annotate each research idea with its publication date, which we use as a cutoff date for literature search during the execution of baselines and ScholarEval.
@@
LLM-based extraction. After identifying 130 papers across the four disciplines of interest satisfying our criteria, we follow a multi-step approach to extract the research idea and its corresponding reviews. (1) Document parsing. We first use GROBID to parse the paper into different sections and employ a heuristic on section names to only keep those related to the background and methodology, while dropping all sections mentioning experiment execution and results. (2) Research idea extraction. We instruct an LLM [^3] We use Claude 4 Sonnet. to extract the research idea based on the parsed documents such that each research idea is composed of three components (Problem, Methods, Experiments). (3) Review rubrics extraction. Finally, we provide an LLM with the full reviews corresponding to the paper along with the extracted research idea and instruct it to extract statements from the reviews pertaining to assessments of the underlying research idea These statements are further classified by the LLM based on their type (strength or weakness), severity (major or minor), and axis (soundness or contribution) to construct our review rubrics.
@@
Expert Validation. We invite 6 subject-matter experts (PhD students, postdocs, and professors) in artificial intelligence, neuroscience, biochemistry, and ecology to validate the quality of each (research idea, review rubrics) pair extracted by the LLM. Namely, the experts are instructed to ensure that the research idea is a faithful representation of the paper at the ideation stage, that the review rubrics do not have mentions to execution, results, or paper presentation, and that the research idea and review rubrics are consistent (i.e. the review only addresses aspects contained in the research idea). Annotators are also asked to verify the correctness of the dimension assignments made by the LLMs. Experts are instructed to make necessary corrections and to discard instances that fall out of their specific area of expertise.
@@
At the end of this process, we obtain 117 validated (research idea, review rubrics) pairs balanced across four disciplines. Figure 3 shows the statistics of ScholarIdeas and an example pair. Further details about ScholarIdeas curation are given in Appendix C.

### v2

@@ Abstract
As AI tools become increasingly common for research ideation, robust evaluation is critical to ensure the validity and usefulness of generated ideas. We introduce ScholarEval, a retrieval-augmented evaluation framework that assesses research ideas based on two fundamental criteria: soundness—the empirical validity of proposed methods based on existing literature, and contribution—the degree of advancement made by the idea across different dimensions relative to prior research. To evaluate ScholarEval, we introduce ScholarIdeas, the first expert-annotated dataset of multi-domain research ideas and reviews, comprised of 117 ideas across four disciplines: artificial intelligence, neuroscience, biochemistry, and ecology. Our evaluation shows that ScholarEval achieves significantly higher coverage of points mentioned in the human expert annotated rubrics in ScholarIdeas compared to all baselines. Furthermore, ScholarEval is consistently preferred over the strong baseline OpenAI Deep Research, a reasoning and search-enabled agentic system, in terms of evaluation actionability, depth, and evidence support. Our large-scale user study also shows that ScholarEval significantly outperforms OpenAI Deep Research in literature engagement, idea refinement, and usefulness. We openly release our code, dataset, and ScholarEval tool for the community to use and build on. [^1] Code and data can be found at https://github.com/skai-research/ScholarEval
ScholarEval: Research Idea Evaluation
Grounded in Literature
Hanane Nour Moussa1222Co-leads. All authors’ contributions are detailed in the Contribution section., Patrick Queiroz Da Silva1222Co-leads. All authors’ contributions are detailed in the Contribution section., Daniel Adu-Ampratwum2, Alyson East3, Zitong Lu4, Nikki Puccetti5, Mingyi Xue6, Huan Sun1, Bodhisattwa Prasad Majumder7, Sachin Kumar1 1Department of Computer Science and Engineering, The Ohio State University 2Division of Medicinal Chemistry and Pharmacognosy, The Ohio State University 3Department of Wildlife, Fisheries, and Conservation Biology, The University of Maine 4McGovern Institute for Brain Research, Massachusetts Institute of Technology 5Center for Cognitive and Behavioral Brain Imaging, The Ohio State University 6Department of Chemistry, University of Wisconsin-Madison 7Allen Institute for Artificial Intelligence
@@ 1 Introduction
Figure 1: Top: Given a research idea ScholarEval generates a literature-grounded evaluation based on soundness and contribution. Bottom: To evaluate ScholarEval, we measure the degree of coverage of expert-annotated review rubrics in ScholarIdeas. The final coverage score is the average over all rubrics.
@@
Despite this need, automatic research idea evaluation remains an underexplored area. Some works narrowly frame it as a prediction task: deciding which idea among a pair would lead to better results on predefined benchmarks (Wen et al., 2025). Others focus on one-dimensional approaches to idea evaluation (e.g., novelty) (Afzal et al., 2025; Shahid et al., 2025), or target only specific sub-disciplines (e.g., AI) (Si et al., 2025b). Importantly, these systems either only generate scores or are limited to sparse feedback such as short rationale statements (Feng et al., 2025; Wen et al., 2025; Shahid et al., 2025). However, to create AI co-scientists that can generate and refine research ideas, giving dense, actionable, and multifaceted feedback is crucial (Wu et al., 2023; Cao et al., 2024). To the best of our knowledge, no existing work addresses the challenge of comprehensive research idea evaluation across disciplines within a framework that provides detailed and actionable feedback for idea refinement.
We introduce ScholarEval (Figure 1), a research idea evaluation system grounded in the most recent literature. ScholarEval evaluates research ideas based on two fundamental criteria: soundness and contribution. (1) Soundness refers to the empirical validity of each proposed method in the research plan, assessed by systematically examining whether similar applications of this method in existing literature have demonstrated success or failure. (2) Contribution refers to the degree of advancement a research idea offers across different dimensions–e.g., its proposed methodology, data, evaluation approaches, and conceptual framework–relative to existing literature. The rationale for dimension-based evaluation is that novelty is multi-faceted by nature, and an idea can be considered novel relative to certain aspects of prior work, rather than being categorically novel or not (Rubaiat et al., 2025; Radensky et al., 2025). Given a research idea detailing the problem, proposed methodology, and planned experiments, ScholarEval employs a multi-stage pipeline that first generates targeted search queries to retrieve a large volume of related literature from Semantic Scholar (Kinney et al., 2025) (Figure 2). It then extracts key information from the retrieved literature to assess the research plan’s soundness and contribution. Finally, it synthesizes detailed feedback supported by relevant citations.
@@
To evaluate ScholarEval, we construct ScholarIdeas, a multi-disciplinary dataset of 117 research ideas and their corresponding reviews validated by subject-matter experts across four disciplines: artificial intelligence, biochemistry, neuroscience, and ecology. As showcased in Figure 1, reviews in ScholarIdeas are composed of multiple rubrics, each focusing on a specific point that the evaluation should address, for 1076 rubrics in total. We develop a multi-faceted automatic evaluation framework to assess ScholarEval against strong baselines, namely state-of-the-art LLMs and deep research systems. Our results show that ScholarEval achieves greater coverage of the expert-annotated review rubrics in ScholarIdeas, significantly outperforming all baselines and surpassing OpenAI Deep Research by over 20% relative improvement. Our results also demonstrate that ScholarEval is consistently preferred over OpenAI Deep Research in terms of evidence support, depth, and actionability.
@@
Our human study involving 18 experts and 46 evaluations further supports the real world usefulness of ScholarEval across our four target disciplines. ScholarEval outperforms deep research by a significant margin in metrics tied to our core contributions: literature engagement, citation use, feedback validity, idea refinement, relevant evaluation aspects, and overall usefulness. Our work makes the following major contributions:
@@
  • ScholarEval, a literature-grounded framework that comprehensively evaluates research ideas based on their soundness and contribution with actionable feedback. We will openly release the ScholarEval tool and user interface.
@@ 2 ScholarEval
Task Formulation. Given a research idea $I$, the task is to find papers $P=\{p_{1},p_{2},p_{3},\ldots\}$ that are highly relevant to $I$ and synthesize their findings in the context of $I$ to generate a comprehensive evaluation $\mathcal{E}=(S,C)$ covering its soundness and contribution. The evaluation is accompanied by citations, ensuring that all claims are supported by evidence from the literature.
Overview of ScholarEval. As shown in Figure 2, ScholarEval is composed of two main modules: Soundness and Contribution. We present both their workflows in §2.1 and §2.2 and provide further details in Appendix B.
@@ 2.2 Contribution Evaluation
Dimension Extraction. The initial step of contribution evaluation consists of extracting the dimensions along which the research idea is making contributions to the field. Dimensions represent the facets of the idea’s potential contributions that are specific and comparable across related literature (e.g., system design, data collection, evaluation methodology, etc.). Instead of imposing pre-defined dimensions as in Radensky et al. (2025), we use LLM extraction to ensure flexibility based on the nature of the research idea (Rubaiat et al., 2025). Specifically, given the research idea $I$, we use $\mathcal{M}$ to extract dimensions $D=\{d_{1},d_{2},\ldots,d_{l}\}$. Examples of dimensions include tool or system design, conceptual framework, evaluation methodology, etc. Each $d_{i}$ also includes the reasoning for how the idea makes contributions along that dimension. These statements will provide important context to ground the query generation in the subsequent step.
@@
Paper Discovery. In this step, we conduct a broad search over the literature to identify relevant papers to compare $I$ against. However, unlike the soundness module, which requires searching paper content for methodological details, contribution evaluation can be performed using only abstracts, since a paper’s main contributions are typically highlighted there. This also allows us to cast a wider net and gather a broad set of related papers. Such breadth is essential for contribution evaluation, as determining truly novel contributions requires an exhaustive view of the literature. The paper discovery thus consists of the following steps: (1) For each extracted dimension, $\mathcal{M}$ generates queries to search for relevant papers using Semantic Scholar paper search and retrieve their abstracts. (2) $\mathcal{M}$ assesses similarity in contributions of each candidate paper abstract relative to the research idea $I$ and assigns a score on a scale from 1 to 5. Papers that are deemed highly relevant (i.e., score $\geq 3$) are then used as seeds for the paper augmentation stage. (3) Paper augmentation leverages the Semantic Scholar Recommendations API to find similar papers, and we additionally extract the publications cited by each seed paper. (4) This augmented list of candidate papers then undergoes another stage of relevance assessment by $\mathcal{M}$. However, due to the typically large volume of papers at this stage, we first filter this list to the top $n$ papers based on semantic embedding [^2] We use Titan Text Embedding v2 (Amazon Web Services, 2025) similarity between abstracts and $I$ before forwarding to $\mathcal{M}$.
@@
Pairwise Comparison. Once the final set of papers $P_{D}=\{p_{1},p_{2},\ldots,p_{m}\}$ is identified, we conduct a series of pairwise comparisons to uncover how these papers’ contributions compare to those proposed in $I$. Specifically, we prompt $\mathcal{M}$ to compare each paper’s abstract to $I$ along each dimension $d_{i}\in D$. This comparison produces a granular view of the areas in which $I$ is making novel advancements and those in which it is lacking in novelty, compared to existing work.
@@
Contribution Review Synthesis. The final stage of contribution evaluation consists of synthesizing dimension-level contribution assessments. For each $d_{i}\in D$, $\mathcal{M}$ uses the results of the pairwise comparisons to synthesize an evaluation composed of three sections: (1) Strengths: the novel contributions that $I$ makes along this dimension compared to prior work. (2) Weaknesses: the areas in which the contributions of $I$ are lacking or limited compared to prior work. (3) Suggestions: actionable recommendations to improve the novelty of $I$ along this dimension.
@@ 3 ScholarIdeas: A dataset for research idea evaluation
To this end, we curate ScholarIdeas, a dataset containing research ideas and reviews covering four disciplines: artificial intelligence, neuroscience, biochemistry, and ecology. Due to a lack of sources with pre-execution research ideas, we employ a semi-automatic pipeline to retrospectively extract ideas from existing papers, where each example undergoes rigorous evaluation by subject-matter experts (§3.1). Our primary motivation for developing ScholarEval is to evaluate AI-generated ideas. However, collecting detailed expert reviews for AI-generated ideas at scale is prohibitively expensive and practically infeasible. Our dataset of existing human-written ideas with expert reviews serves as a reliable proxy.
The evaluation of long-form responses, such as the ones generated by ScholarEval, is also inherently challenging. While other scientific tasks admit easily verifiable success criteria (e.g., code generation (Jansen et al., 2025; Chen et al., 2025; Li et al., 2025b)), such evaluation metrics are not readily available for our use case. We develop a multi-faceted automatic evaluation pipeline detailed in §3.2. We further corroborate our automatic evaluation with a large-scale human evaluation (§5).
@@ 3.1 Data Curation and Annotation
Data selection. We manually select papers and reviews from two sources: OpenReview (ICLR 2025) for AI-related research ideas and eLife for life sciences. Both sources include high-quality reviews by multiple reviewers for all submitted papers, in contrast to many other sources that do not openly release reviews or restrict them to those of accepted manuscripts. To ensure the suitability of the reviews to be used as ground-truth, we only select papers satisfying the following criteria: (1) The paper must be reviewed by at least two reviewers. (2) All reviews of the paper are in agreement and offer a general consensus about the quality of the work. (3) The reviews mention criticism about the underlying research idea and not exclusively about obtained results or other details known post-execution. For each of the sources we use, we only retrieve the first version of the submission prior to any revisions, along with the first round of reviews. We also annotate each idea with its publication date, which we use as a cutoff date for literature search during ScholarEval and baseline execution.
@@
LLM-based extraction. After identifying 130 papers across the four disciplines of interest satisfying our criteria, we follow a multi-step approach to extract the research idea and its corresponding reviews. (1) Document parsing. We first use GROBID to parse the paper into different sections and employ a heuristic on section names to only keep those related to the background and methodology, while dropping all sections mentioning experiment execution and results. (2) Research idea extraction. We instruct an LLM [^3] We use Claude 4 Sonnet to extract the research idea based on the parsed documents such that each research idea is composed of three components (Problem, Methods, Experiments). (3) Review rubrics extraction. Finally, we provide an LLM with the full reviews corresponding to the paper along with the extracted research idea and instruct it to extract statements from the reviews pertaining to assessments of the underlying research idea. These statements are further classified by the LLM based on their type (strength or weakness), severity (major or minor), and axis (soundness or contribution) to construct our review rubrics.
@@
Expert Validation. We invite 6 subject-matter experts (PhD students, postdocs, and professors) in artificial intelligence, neuroscience, biochemistry, and ecology to validate the quality of each (research idea, review rubrics) pair extracted by the LLM. Namely, the experts are instructed to ensure that the research idea is a faithful representation of the paper at the ideation stage, that the review rubrics do not mention execution, results, or paper presentation, and that the research idea and review rubrics are consistent (i.e. the review only addresses aspects contained in the research idea). Annotators are also asked to verify the correctness of the dimension assignments made by the LLMs. They are instructed to make necessary corrections and to discard instances not in their area of expertise.
@@
At the end, we obtain 117 validated (research idea, review rubrics) pairs balanced across four disciplines. Figure 3 shows the statistics of ScholarIdeas and an example pair. Further details about ScholarIdeas curation are given in Appendix D.